OpenStudy (anonymous):

The excursion boat on the river takes 2½ hours to make the trip to a point 12 miles upstream and to return. If the rate at which the boat travels in still water is 5 times the rate of the river current, what is the rate of the current?

OpenStudy (anonymous):

x= Current rate Boat in Still water =B= (2.5)(B-x)=24 B=5x

OpenStudy (anonymous):

2.5B-2.5x=24 B=5x Using Subsitution 2.5(5x)-2.5(x)=24 Solve for x

OpenStudy (anonymous):

2.4?

OpenStudy (anonymous):

Yes